# Leveraging Auxiliary Task Relevance for Enhanced Industrial Fault Diagnosis through Curriculum Meta-learning

### v1

@@ IV Conclusion and Future Work
In this paper, we develop the Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, a novel approach in the realm of fault diagnosis for smart manufacturing designed to address the issue of data scarcity. RT-ACM innovatively utilizes auxiliary tasks to enrich the primary fault diagnosis tasks, adhering to the principle of “paying more attention to more relevant knowledge” by focusing on the most relevant knowledge and adopting an “easy-to-hard” curriculum learning strategy. Our comprehensive evaluations across three real-world datasets confirm the effectiveness and superiority of the RT-ACM framework. The experiments not only highlight the efficacy of the framework but also validate the architectural choices, underscoring the potential of combining curriculum strategy with related task aware meta-learning for enhanced fault diagnosis. In future work, we will explore new strategies that can learn both temporal and spatial representations of auxiliary tasks when transferring knowledge to the target task.

### v2

@@ IV Conclusion and Future Work
In this paper, we introduced a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, a novel approach in the realm of fault diagnosis for smart manufacturing designed to address the issue of data scarcity. RT-ACM introduces an innovative approach by employing related task-aware meta-learning to enhance primary fault diagnosis tasks. It adheres to the principle of “paying more attention on relevant knowledge” by prioritizing the most pertinent information and following an “easy-to-hard” curriculum strategy. Our comprehensive evaluations across two real-world datasets confirm the effectiveness and superiority of the RT-ACM framework. The experiments not only highlight the efficacy of the framework but also validate the architectural choices, underscoring the potential of combining curriculum strategy with related task aware meta-learning for enhanced fault diagnosis. In future work, we will explore new strategies that can learn both temporal and spatial representations of auxiliary tasks when transferring knowledge to the target task.